# Action-OOD: An End-to-End Skeleton-Based Model for Robust Out-of-Distribution Human Action Detection

### v1

@@ 3.1 GCN-based feature extractor
Since many GCN based models have achieved great success in action recognition tasks, in this work we use existing models that perform best in ID data classification to extract embeddings from human skeleton data. It is an important step for a GCN based model to construct the graph topology in advance. A topology can bring great convenience to the spatial learning of the model while incorporating prior information thus improving the physical consistency of the model. Following the previous work by (Lee et al., 2023), we define three types of links for joint connections, while keeping the number of nodes unchanged (see Fig. 2).
Figure 2: Description of the graph construction method. Taking the Kinetics dataset as an example, the left part is the visualized body joints and their physical connections. The right part is the transformed balance tree based on PC edges. The orange edges represent the self-loop of joints. Green edges represent fully connected edges within each centripetal joint subset.
Let us define $G=\left\{{g_{pc},g_{sl},g_{fc}}\right\}$ as the three types of graph subsets. The first type of subset $g_{pc}$ is physical connections, which are represented by blue lines between joints in Fig. 2. It represents the natural physical relationships in skeleton data. The second type $g_{sl}$ is the self-loop, where nodes are connected to themselves. By incorporating self-loops, nodes are enabled to consider their own features during message passing in addition to relying on neighboring nodes’ information, which helps stabilize the training process. Additionally, self-loops can be regarded as a regularization mechanism to prevent over-fitting, making them a commonly used method in graph neural network construction(Kipf and Welling, 2017).
The third type of subset is a hierarchical fully connected joints $g_{fc}$. We employ a balanced tree approach to re-represent the graph structure (as illustrated on the right side of Fig. 2). Fully connected edges are constructed between adjacent nodes across two levels of the balanced tree, based on the hierarchical level of the nodes. This approach considers that the characteristics of actions may correlate with features across various levels of the body’s hierarchical structure. Additionally, it expands the model’s receptive field spatially in the feature space, enhancing the model’s generalization ability. This approach has been widely proven to be effective in previous experiments. After that, for the input skeleton data $X$, the output feature maps $F$ as:
$F=f^{HD-GCN}(X,G).$ (4)
@@ 3.2 Feature activation shaping and fusion
Following the previous work (Djurisic and et al., 2023), we utilize activation shaping (ASH) operation to filter out useless information in high-dimensional HD-GCN captured feature maps $F$, which can bring benefits to other downstream tasks (e.g. OOD detection). However, the difference is that we utilize it in the training stage to help the model itself distinguish OOD samples.
@@
There are three types of ASH strategies: ASH-P, ASH-B, and ASH-S. As shown in Fig. 3, for the input feature maps $F$, ASH-P only set values less than $t$ into zero, ASH-B reset the non-zeros by binarization, and ASH-S set them with scaling. Details of these three strategies can be found in supplementary material. The difference among them lies in the normalization strategies after pruning. In the experimental part, we show the results of all these three strategies in training. After that, we retain both the output feature maps and the pruning features:
$\hat{F}=\left[F,\text{ASH}(F)\right],$ (5)
where ASH($F$) represents the results of any activation strategy acting on the feature maps $F$, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the process to concatenate two vectors into a single one. The concatenated features undergo two main operations in the feature fusion process: one SE layer (Hu et al., 2018) primarily aimed at enhancing the model’s sensitivity to channel features:
@@ 3.3 OOD detection using logits
Here Eqn. 8 shows the classification operation after data fusion and $W_{fc}^{\prime}$ are the learnable weights. Then we get the logits ready for the downstream OOD detection task. The dimension of logits is equivalent to the total number of seen classes plus one unseen class. The detailed process of OOD detection is illustrated in Fig. 4.
@@
For the input logits after the classifier, we measure the energy score of each sample using Eqn. 3: If the score is smaller than $\tau$, we will recognize it as ‘unseen’ for OOD and set the element value representing the probability that the sample belongs to the unknown category to 1. Otherwise, the probability of unseen will remain zero and others will be the value between 0 to 1. The value of parameter $\tau$ is given based on the top 10% of energy scores on the training set. The assumption behind this is that the energy score of OOD after training is less than the energy score of 90% ID data. After that, we can get a vector $\hat{Y}$ that represents the probability that the sample belongs to each class.
Algorithm 1 Action-OOD Detection Algorithm
@@
1:   $F=f^{HD-GCN}(X,G)$triangleright extract feature maps
@@ 3.4 Energy based loss function
The first part is the cross-entropy part, where $P(logits_{in})$ is the expected probability output, and $f(\cdot)$ represents the model output on in-distribution training data. The second part is energy-based, where $\mathcal{D}_{\text{in}}^{\text{train}}$ is the in-distribution training data, $E\left(logits_{in}\right)$ represents the energy score of in-distribution data, $m_{in}$ is the margin parameter which is set before can help the ID energy score distribution reduce the variance and close to this point, $\alpha$ is the tuning parameter that adjusts the proportion of energy part. With the help of this learning objective, the variance of ID data will be reduced. We can also ensure the model understand the energy score distribution of ID data, thus indirectly having the ability to distinguish ID and OOD samples. The overall process can be found in Algorithm 1.
@@ 4 Experiments
To demonstrate the human action detection accuracy of our proposed model Action-OOD, a series of experiments are conducted. In this section, we elucidate the intricate details of the experimental design and present a thorough analysis of the experimental results. First, we introduce three datasets and describe their data pre-processing methods. Subsequently, we compare our work with other existing OOD detection methods and identify the proposed model’s strengths and weaknesses. Finally, we design three ablation studies to examine the effectiveness of individual modules.
@@ 4.1 Dataset description
    NTU-RGB+D 60 (Shahroudy et al., 2016) & 120 (Liu et al., 2019)  [^1] https://rose1.ntu.edu.sg/dataset/actionRecognition/: The NTU-RGB+D 60 dataset (NTU60) comprising 60 motion categories and 56,880 video samples, is extensively employed in diverse action recognition tasks. The NTU-RGB+D 120 dataset (NTU120) is an extension of the NTU-RGB+D 60 dataset, expanding the total number of human action categories to 120. Both datasets encompass 5 data modalities: 3D skeletons (body joints), masked depth maps, full-depth maps, RGB videos, and IR data. In this study, we exclusively utilize the 3D skeleton sequence data. The 3D skeleton data includes the 3D coordinates (X, Y, Z) of 25 body joints per frame and a maximum of 2 subjects. The actions in these datasets are categorized into three main groups: daily actions, interactive actions, and medical conditions. Five classes are randomly selected as unseen from the NTU60 dataset and ten classes from the NTU120 dataset. Note that these unseen categories encompass samples from all three categories.
@@
    Kinetics human action dataset (Kinetics 400) (Kay et al., 2017): It is by far the largest unconstrained action recognition dataset constructed by DeepMind. It covers 400 action classes from human-object interactions to human-human interactions and some other complex interactions from YouTube videos. Since the Kinetics dataset provides only raw video clips without skeleton data, we adopt the skeleton-based action data processed by previous work STGCN (Yan et al., 2018)  [^2] https://github.com/yysijie/st-gcn. The processed data sequences contain 18 joints and 2 subjects. We randomly select 33 action classes marked as ‘unseen’ for OOD samples, ensuring the same ratio as used in NTU60 and NTU120.
@@ 4.2.1 Experiment settings
    Implementation details: We adopt HD-GCN (Lee et al., 2023) as the backbone to extract Graph-structured skeleton data as shown in Fig. 1. The batch size of data is 64. The SGD optimizer is employed with a Nesterov momentum of 0.9 and a weight decay of 0.0004. The learning rate is 0.001 for NTU60 and NTU120 and 0.01 for Kinetics400. The number of learning epochs is set to 100 for NTU60 and NTU120, and 150 for Kinetics400, with a warm-up strategy (He et al., 2016) applied to the first five epochs for more stable learning. All our experiments are conducted on NVIDIA GeForce RTX 3090, CUDA 11.6 + PyTorch 1.12.1.
@@
    Evaluation metric: As depicted in Fig. 4, we assess both the accuracy of detecting unseen samples and classifying seen samples. All samples under evaluation first undergo a threshold-based determination to discern whether their labels are unseen. If classified as ‘unseen’, they are marked accordingly; otherwise, they are predicted to belong to known categories. Regarding the unseen detection task, it can be viewed as a binary classification problem. We evaluate its accuracy using Detection Error (Liang et al., 2019), AUROC (Davis and Goadrich, 2006), and FPR at 95% TPR (Liang et al., 2019), following the methodology outlined in (Zhao et al., 2023). For action recognition tasks involving seen data, we employ Top-1 Accuracy (Lee et al., 2023) for quantitative evaluation. The details of these evaluation metrics are outlined as follows:
@@
      Detection Error (Error) (Liang et al., 2019): It measures the misclassification probability when the True Positive Rate (TPR) is 95%. The definition of an Error is:
@@
      AUROC (Davis and Goadrich, 2006): It stands for the Area Under the Receiver Operating Characteristic (ROC) curve, which depicts the relation between TPR and FPR and ranges from 0 to 1 (the larger the value, the higher the accuracy). AUROC provides a single scalar value that summarizes the overall performance of a binary classifier across all possible thresholds.
@@
      FPR at 95% TPR (FPR95) (Liang et al., 2019): It refers to the rate of false positive predictions when the true positive rate reaches 95% in a classification task.
@@ Methods
    4. 4)
      Top-1 Accuracy (ACC) (Lee et al., 2023): It measures the proportion of correctly predicted instances among all instances, where the predicted class label exactly matches the ground truth label for each instance. It is defined as:
      $ACC=\frac{hit}{N}\times 100\%,$ (11)
    5. 5)
  • 3.
4.2.2 Baselines
Since most OOD detection models are post-hoc methods, we retain basic feature extraction parts and change different post-hoc methods to achieve OOD detection. We compare Action-OOD with four main methods: one score measuring method (Maximum Softmax Probability, MSP (Hendrycks and Gimpel, 2017)) and three different feature activation methods – ReAct (Rectified Activation) (Sun et al., 2021), DICE (Directed Sparisification) (Sun and Li, 2022), DICE+ReAct, ASH (Activation Shaping) (Djurisic and et al., 2023). Details are shown below:
  • 1.
    MSP (Maximum Softmax Probability): (Hendrycks and Gimpel, 2017) claims that out-of-distribution examples tend to have lower maximum softmax probabilities than known samples. This scoring-based method has shown wide success in the computer vision area.
  • 2.
    ReAct (Rectified Activation) (Sun et al., 2021): A post-hoc method proposed to reduce the overconfidence problem of the model on OOD data by truncating activations on the penultimate layer of a network to limit the effect of noise.
  • 3.
    DICE (Directed Sparisification) (Sun and Li, 2022): Another post-hoc method, based on the idea of denoising useless information, sorts the weights of its contribution (i.e., weight × activation). It claims that only a subset of units contributes to the in-distribution prediction results. However, it will result in non-negligible noise signals when measuring out-of-distribution data.
  • 4.
    DICE + ReAct: Following previous work (Djurisic and et al., 2023), we construct a baseline that filters out noise signals using both DICE and ReAct methods.
  • 5.
    ASH (Activation Shaping) (Djurisic and et al., 2023): It consists of three different mechanisms to conduct pruning-based feature activation shaping, including ASH-P, ASH-S, and ASH-B.
4.2.3 Ablation study
To verify the effectiveness, we design ablation studies on the following components of Action-OOD.
  • 1.
    GCN: This is designed to measure the effect of the feature fusion block of Action-OOD. As previously mentioned, Action-OOD has two parts of features: GCN-based features, which excel in classifying in-distribution (ID) samples, and ASH-based features, which aim to reduce overconfidence in predictions for OOD samples. This experiment can prove the improvement effect of ASH operation on the model OOD recognition task.
  • 2.
  • 3.
@@ 4.2.3 Ablation study
ID NTU60 NTU120 Kinetics400 Average
@@
4.3 Results and discussion
4.3.1 Performance comparison against baselines
We evaluate the proposed Action-OOD method against baselines across three datasets for two tasks: (a) to distinguish OOD samples from an ID and OOD mixture dataset, and (b) to evaluate whether ID samples will be misclassified as OOD merely based on ID data. We also provide the average results for all models involved across the three datasets.
It can be observed that compared to the first two datasets, the overall prediction accuracy on Kinetics is not satisfactory. Additionally, the direct classification results on the ID dataset are relatively poor. By comparing the incorrectly predicted samples on the Kinetics data set, we discover that some ID samples are misclassified as OOD samples, which means that the model lacks confidence in the ID samples. This also shows to some extent that the data itself may lack expressive power, which prevents the model itself from learning more distinguishing features. One possible reason could be attributed to the graph construction as shown in Fig. 2. Firstly, the Kinetics dataset itself includes a large number of actions involving interactions between people and objects, making it difficult to distinguish using skeleton data only. Secondly, through our visualization analysis, there are certain issues in extracting Kinetics nodes in the STGCN (Yan et al., 2018). For instance, nodes 15 and 16 in the data represent the left and right eyes respectively. However, these nodes may not provide much information about the actions. Moreover, while the dataset models the left and right hips separately, it lacks crucial information about the key body waist nodes, which are essential for understanding the type of limb movement.

### v2

@@ 3.1 GCN-based feature extractor
Since many GCN-based models have achieved great success in action recognition tasks, in this work we use existing models that perform best in ID data classification to extract embeddings from human skeleton data. It is an important step for a GCN-based model to construct the graph topology in advance. A topology can bring great convenience to the spatial learning of the model while incorporating prior information thus improving the physical consistency of the model. Following the previous work by [23], we define three types of edges for joint connections, while keeping the number of nodes unchanged (see Figure 2).
Figure 2: Description of the graph construction method. (a) Physically connected edges. The nodes and edges marked in each subset correspond to the nodes and edges of the two adjacent layers of nodes in the rightmost tree structure. (b) Self-loop edges (green arrow pointing to the joints themselves). We add self-loops of each joints to help the model take into account the characteristics of the joints themselves when conducting graph convolution. (c) Fully connected inter-hierarchy edges. We assume that the inter-hierarchy relationships in each subsets can help to distinguish different actions. Thus, we add fully-connected edges in each subsets as shown in yellow lines.
Let us define $G=\left\{{g_{pc},g_{sl},g_{fc}}\right\}$ as the three types of graph. The first type of graph $g_{pc}$ is physical connected edges, which are represented by blue lines between joints in Figure 2 (a). It represents the natural physical relationships in skeleton data. The second type $g_{sl}$ is the self-loop, where nodes are connected to themselves. By incorporating self-loops, nodes are enabled to consider their own features during message passing in addition to relying on neighboring nodes’ information, which helps stabilize the training process. Additionally, self-loops can be regarded as a regularization mechanism to prevent over-fitting, making them a commonly used method in graph neural network construction[31].
The third type of graph is a hierarchical fully connected graph $g_{fc}$. We employ a balanced tree approach to reconstruct the graph structure (as illustrated on the right side of Figure 2). Fully connected edges are constructed between adjacent nodes across two levels of the balanced tree, based on the hierarchical level of the nodes. This approach considers that the characteristics of actions may correlate with features across various levels of the body’s hierarchical structure. Moreover, it expands the model’s receptive field spatially in the feature space, enhancing the model’s generalization ability. This approach has been widely proven to be effective in previous experiments. After that, for the input skeleton joints data $X$, graph subsets $G$ and temporal sequence $T$, the output feature maps $F$ are obtained as follows:
$F=f^{HD-GCN}(X,G,T).$ (4)
@@ 3.2 Feature activation shaping and fusion
Following the previous work [14], we utilize activation shaping (ASH) operation to filter out useless information in high-dimensional HD-GCN captured feature maps $F$, which can bring benefits to other downstream tasks (e.g. OOD detection). However, the difference is that we utilize it in the training stage to help the model itself distinguish OOD samples.
There are three types of ASH strategies: ASH-P, ASH-B, and ASH-S. As shown in Figure 3, for the input feature maps $F$, ASH-P only set values less than $t$ into zero, ASH-B reset the non-zeros by binarization, and ASH-S set them with scaling. Details of these three strategies can be found in A. The difference among them lies in the normalization strategies after pruning. In the experimental part, we show the results of all these three strategies in training. After that, we retain both the output feature maps and the pruning features:
$\hat{F}=\left[F,\text{ASH}(F)\right],$ (5)
@@
where ASH($F$) represents the results of any activation strategy acting on the feature maps $F$, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the process to concatenate two vectors into a single one. The concatenated features undergo two main operations in the feature fusion process: one SE layer [32] primarily aimed at enhancing the model’s sensitivity to channel features:
@@ 3.3 OOD detection using logits
Here Eqn. 8 shows the classification operation after data fusion, and $W_{fc}^{\prime}$ is the learnable weight. Then we get the logits ready for the downstream OOD detection task. The logits’ dimension equals the number of seen classes plus one unseen class. The detailed process of OOD detection is illustrated in Figure 4.
@@
For the input logits after the classifier, we measure the energy score of each sample using Eqn. 3: If the score is smaller than $\tau$, we will recognize it as ‘unseen’ for OOD and set the element value representing the probability that the sample belongs to the unknown category to 1. Otherwise, the probability of unseen will remain 0 and others (represent specific seen classes) will be the value between 0 to 1. The output of this stage will be multi-classes. The value of parameter $\tau$ is given based on the top 10% of energy scores on the training set. The assumption behind this is that the energy score of OOD after training is less than the energy score of 90% ID data. After that, we can get a vector $\hat{Y}$ that represents the probability that the sample belongs to each class.
Algorithm 1 Skeleton-OOD Detection Algorithm
@@
1:   $F=f^{HD-GCN}(X,G,T)$triangleright extract feature maps
@@ 3.4 Energy based loss function
The first part is the cross-entropy part, where $P(logits_{in})$ is the expected probability output, and $f(\cdot)$ represents the model output on in-distribution training data. The second part is energy-based, where $\mathcal{D}_{\text{in}}^{\text{train}}$ is the in-distribution training data, $E\left(logits_{in}\right)$ represents the energy score of in-distribution data, $m_{in}$ is the margin parameter which is set before can help the ID energy score distribution reduce the variance and close to this point, $\alpha$ is the tuning parameter that adjusts the proportion of energy part. With the help of this learning objective, the variance of ID data will be reduced. We can also ensure the model understands the energy score distribution of ID data, thus indirectly having the ability to distinguish ID and OOD samples. The overall process can be found in Algorithm 1.
@@ 4 Experiments
To demonstrate the human action detection accuracy of our proposed model Skeleton-OOD, a series of experiments are conducted. In this section, we elucidate the intricate details of the experimental design and present a thorough analysis of the experimental results. First, we introduce three datasets and describe their data pre-processing methods. Subsequently, we compare our work with other existing OOD detection methods and identify the proposed model’s strengths and weaknesses. Finally, we design three ablation studies to examine the effectiveness of individual modules.
@@ 4.1 Dataset description
    NTU-RGB+D 60 [34] & 120 [35]  [^1] https://rose1.ntu.edu.sg/dataset/actionRecognition/: The NTU-RGB+D 60 dataset (NTU60), comprising 60 motion categories and 56,880 video samples, is extensively employed in diverse action recognition tasks. The NTU-RGB+D 120 dataset (NTU120) is an extension of the NTU-RGB+D 60 dataset, expanding the total number of human action categories to 120. Both datasets encompass five data modalities: 3D skeletons (body joints), masked depth maps, full-depth maps, RGB videos, and IR data. In this study, we exclusively utilize the 3D skeleton sequence data. The 3D skeleton data includes the 3D coordinates (X, Y, Z) of 25 body joints per frame for a maximum of two subjects. The actions in these datasets are categorized into three main groups: daily actions, interactive actions, and medical conditions. Five classes are randomly selected as unseen from the NTU60 dataset and ten classes from the NTU120 dataset. Note that these unseen categories encompass samples from each of the three groups.
@@
    Kinetics human action dataset (Kinetics 400) [36]: It is, by far, the largest unconstrained action recognition dataset created by DeepMind. It covers 400 action classes, ranging from human-object interactions to human-human interactions, and also includes other complex interactions from YouTube videos. Since the Kinetics dataset provides only raw video clips without skeleton data, we adopt the skeleton-based action data processed by previous work STGCN  [^2] https://github.com/yysijie/st-gcn [1]. The processed data sequences contain 18 joints and 2 subjects. We randomly select 33 action classes marked as ‘unseen’ for OOD samples, ensuring the same ratio as used in the NTU60 and NTU120 datasets.
@@ 4.2.1 Experiment settings
    Implementation details: We adopt HD-GCN [23] as the backbone to extract graph-structured skeleton data as shown in Figure 1. The batch size of data is 64. The SGD optimizer is employed with a Nesterov momentum of 0.9 and a weight decay of 0.0004. The learning rate is 0.001 for NTU60 and NTU120 and 0.01 for Kinetics400. The number of learning epochs is set to 100 for NTU60 and NTU120, and 150 for Kinetics400, with a warm-up strategy [37] applied to the first five epochs for more stable learning. All our experiments are conducted on NVIDIA GeForce RTX 3090, CUDA 11.6 + PyTorch 1.12.1.
@@
    Evaluation metric: As depicted in Figure 4, we assess both the accuracy of detecting unseen samples and classifying seen samples. All samples under evaluation first undergo a threshold-based determination to discern whether their labels are unseen. If classified as ‘unseen’, they are marked accordingly; otherwise, they are predicted to belong to known categories. Regarding the unseen detection task, it can be viewed as a binary classification problem. We evaluate its accuracy using Detection Error [38], AUROC [39], and FPR at 95% TPR [38], following the methodology outlined in [40]. For action recognition tasks involving seen data, we employ Top-1 Accuracy [23] for quantitative evaluation. The details of these evaluation metrics are outlined as follows:
@@
      Detection Error (Error) [38]: It measures the misclassification probability when the True Positive Rate (TPR) is 95%. The definition of an Error is:
@@
      AUROC [39]: It stands for the Area Under the Receiver Operating Characteristic (ROC) curve, which depicts the relation between TPR and FPR and ranges from 0 to 1 (the larger the value, the higher the accuracy). AUROC provides a single scalar value that summarizes the overall performance of a binary classifier across all possible thresholds.
@@
      FPR at 95% TPR (FPR95) [38]: It refers to the rate of false positive predictions when the true positive rate reaches 95% in a classification task.
    4. 4)
      Top-1 Accuracy (ACC) [23]: It measures the proportion of correctly predicted instances among all instances, where the predicted class label exactly matches the ground truth label for each instance. It is defined as:
      $ACC=\frac{hit}{N}\times 100\%,$ (11)
    5. 5)
  • 3.
4.2.2 Baselines
Since most OOD detection models are post-hoc methods, we retain basic feature extraction parts and change different post-hoc methods to achieve OOD detection. We compare the proposed Skeleton-OOD with four main methods: one score measuring method MSP (Maximum Softmax Probability) [8] and three different feature activation methods – ReAct (Rectified Activation) [10], DICE (Directed Sparisification) [24], DICE+ReAct, ASH (Activation Shaping) [14]. Details are shown below:
  • 1.
    MSP (Maximum Softmax Probability): [8] claims that out-of-distribution examples tend to have lower maximum softmax probabilities than known samples. This scoring-based method has shown wide success in the computer vision area.
  • 2.
    ReAct (Rectified Activation) [10]: A post-hoc method was proposed to reduce the overconfidence problem of the model on OOD data by truncating activations on the penultimate layer of a network to limit the effect of noise.
  • 3.
    DICE (Directed Sparisification) [24]: Another post-hoc method, based on the idea of denoising useless information, sorts the weights of its contribution (i.e., weight × activation). It claims that only a subset of units contributes to the in-distribution prediction results. However, it will result in non-negligible noise signals when measuring out-of-distribution data.
  • 4.
    DICE + ReAct: Following previous work [14], we construct a baseline that filters out noise signals using both DICE and ReAct methods.
  • 5.
    ASH (Activation Shaping) [14]: It consists of three different mechanisms to conduct pruning-based feature activation shaping, including ASH-P, ASH-S, and ASH-B.
4.2.3 Ablation study
To verify the effectiveness, we design ablation studies on the following components of Skeleton-OOD.
  • 1.
    GCN: This is designed to measure the effect of the feature fusion block of Skeleton-OOD. As previously mentioned, Skeleton-OOD has two parts features: GCN-based features, which excel in classifying in-distribution (ID) samples, and ASH-based features, which aim to reduce overconfidence in predictions for OOD samples. This experiment can prove the improvement effect of ASH operation on the model OOD recognition task.
  • 2.
  • 3.
Except for these, we also explored the results of replacing another backbone as feature extractors and another loss function for training.
4.3 Results and discussion
4.3.1 Performance comparison against baselines
We evaluate the proposed Skeleton-OOD method against baselines across three datasets for two tasks: (a) to distinguish OOD samples from an ID and OOD mixture dataset, and (b) to evaluate whether ID samples will be misclassified as OOD merely based on ID data. We also provide the average results for all models involved across the three datasets.
@@ Methods
At the same time, by comparing the results of the three activation methods of our model horizontally, the performance of the three methods on different data sets has its advantages and disadvantages. As can be seen from Table 2, in the OOD detection task, ASH-P and ASH-B perform better in most indicators. When considering the average performance across all three datasets, the ASH-P activation method demonstrates better dataset transferability and greater stability. From Table 3, our method performs more prominently on the NTU120 and Kinetics data sets on ID data, especially under the activation of ASH-P and ASH-B methods. Judging by the number of top-performing indicators across the three datasets and the average indicator results, ASH-P remains superior. In summary, for real-world OOD detection with Skeleton-OOD, the ASH-P method is recommended for optimal results.
@@
ID sample-only NTU60 NTU120 Kinetics400 Average
@@
In addition, by comparing the results of our method across the three ASH strategies with those of the pure post-hoc ASH strategy, it is evident that incorporating this feature activation operation during the training stage can enhance the model’s OOD detection capability to some extent. For instance, as shown in Table 2, the ASH-P baseline in the fifth row compared to our Ours-ASH-P in the third to last row demonstrates varying degrees of improvement in most indicators across the three datasets. Furthermore, Table 3 illustrates that this approach even slightly enhances the model’s recognition accuracy for ID data.
It can be observed that compared to the first two datasets, the overall prediction accuracy on Kinetics is not satisfactory. Additionally, the direct classification results on the ID dataset are relatively poor. By comparing the incorrectly predicted samples on the Kinetics data set, we discover that some ID samples are misclassified as OOD samples, which means that the model lacks confidence in the ID samples. This also shows to some extent that the data itself may lack expressive power, which prevents the model itself from learning more distinguishing features. One possible reason could be attributed to the graph construction as shown in Figure 2. Firstly, the Kinetics dataset itself includes a large number of actions involving interactions between people and objects, making it difficult to distinguish using skeleton data only. Secondly, through our visualization analysis, there are certain issues in extracting Kinetics nodes in the STGCN [1]. For instance, nodes 15 and 16 in the data represent the left and right eyes respectively. However, these nodes may not provide much information about the actions. Moreover, while the dataset models the left and right hips separately, it lacks crucial information about the key body waist nodes, which are essential for understanding the type of limb movement. In summary, our method outperforms others in certain indicators for some datasets of ID samples and significantly surpasses existing methods in OOD samples. This addresses our primary challenge of enhancing OOD detection capabilities while maintaining ID recognition accuracy.